The demand at Charlottetown’s Queen Elizabeth Hospital for a specialized surgery pack is 70 per week. The lead time from McKesson, its main supplier, is normally distributed, with a mean of five weeks for this product and a standard deviation of two weeks. A 95% weekly service level is desired. Find the ROP.

Note: Round your final answer to the nearest integer.
Answer #1

Reorder point is the quantity level that is reached when the inventory on hand drops to this level. The four determinants of the reorder point quantity are the usage rate, lead time, demand variability, lead time variability and the degree of stock out risk.

ROP= (70x5) + 1.65(70)(2) = 350+ 231 =851